# Score Distillation of Flow Matching Models

### v1

@@ Abstract
Diffusion models achieve high-quality image generation but are limited by slow iterative sampling. Distillation methods alleviate this by enabling one- or few-step generation. Flow matching, originally introduced as a distinct framework, has since been shown to be theoretically equivalent to diffusion under Gaussian assumptions, raising the question of whether distillation techniques such as score distillation transfer directly. We provide a simple derivation—based on Bayes’ rule and conditional expectations—that unifies Gaussian diffusion and flow matching without relying on ODE/SDE formulations. Building on this view, we extend Score identity Distillation (SiD) to pretrained text-to-image flow-matching models, including SANA, SD3-Medium, SD3.5-Medium/Large, and FLUX.1-dev, all with DiT backbones. Experiments show that, with only modest flow-matching- and DiT-specific adjustments, SiD works out of the box across these models, in both data-free and data-aided settings, without requiring teacher finetuning or architectural changes. This provides the first systematic evidence that score distillation applies broadly to text-to-image flow matching models, resolving prior concerns about stability and soundness and unifying acceleration techniques across diffusion- and flow-based generators. We will make the PyTorch implementation publicly available.
Figure 1: Qualitative results of the four-step SiD-DiT generator distilled from SD3.5-Large in a data-free setting ($i.e.$, without access to additional training images).
@@ 1 Introduction
This theoretical equivalence raises a natural question: can diffusion distillation techniques—broadly divided into trajectory and score distillation (Fan et al., 2025), and proven effective for compressing pretrained diffusion models into one- or few-step generators—be directly applied to flow-matching models? Prior work has begun to explore this. The continuous-time consistency model (Lu & Song, 2025) introduced TrigFlow and demonstrated trajectory distillation for pretrained TrigFlow models. Extending this to text-to-image (T2I) generation, Chen et al. (2025) developed SANA-Sprint by reformulating SANA (Xie et al., 2024) from rectified flow into TrigFlow and applying consistency distillation. While effective, this approach requires nontrivial finetuning of rectified-flow checkpoints into TrigFlow counterparts, making it inapplicable to pretrained rectified-flow models without additional adaptation.
Score distillation relaxes the constraint of strictly following the teacher’s sampling trajectory and has shown consistent gains over trajectory-based consistency distillation on diffusion benchmarks such as CIFAR-10 and ImageNet (Zhou et al., 2025c). Yet its applicability to flow-matching T2I models remains unclear. If effective, a further question is whether additional adaptation steps—such as finetuning, as in SANA-Sprint—are necessary. This uncertainty is compounded by a sensitive design space, including noise schedules, loss weighting, network preconditioning (Karras et al., 2022), and architecture. Small changes in these factors can significantly affect performance, as evidenced by methods like sCM, which require careful adaptation during pretraining (Lu & Song, 2025) or finetuning (Chen et al., 2025). Concerns about stability further complicate matters: consistency distillation was favored in SANA-Sprint partly due to instability observed in Distribution Matching Distillation (DMD) (Yin et al., 2024c; a). However, it remains unclear whether this instability is unique to DMD’s KL-based formulation or reflects broader issues in score distillation, which can also be defined with divergences such as Fisher divergence (Zhou et al., 2024) or $f$-divergences (Xu et al., 2025). Huang et al. (2024) argue flow matching does not explicitly model probability density, raising doubts about the soundness of applying distribution-divergence-based objectives directly.
@@
To validate this view, we adopt the few-step Score identity Distillation (SiD) framework (Zhou et al., 2025a), previously shown effective for diffusion models such as SD1.5 and SDXL with U-Net backbones. Here, we extend SiD to pretrained flow-matching models with Diffusion Transformer (DiT) (Peebles & Xie, 2023) backbones, including SANA (Xie et al., 2024; Chen et al., 2025), SD3-Medium, SD3.5-Medium, SD3.5-Large, and FLUX.1-dev (Labs, 2024), spanning 0.6B–12B parameters (2.4–48 GB in fp32).
We demonstrate that SiD, when appropriately configured with automatic mixed precision (AMP) or pure bfloat16 (BF16), operates out of the box in a data-free setting, requiring no images beyond the teacher. In a data-aided setting with additional high-quality text–image pairs, SiD achieves further improvements by incorporating adversarial learning via spatial pooling of discriminator features from a suitable DiT layer, without introducing extra parameters.
We provide a review of related work in Appendix B. Our code will be released to the public. Importantly, a single codebase and hyperparameter configuration suffice across all T2I flow-matching models, underscoring the robustness and applicability of the SiD-DiT framework.
@@ 2.2 Equivalence of Diffusion and Flow-Matching Objectives and Variants
TrigFlow. In TrigFlow (Lu & Song, 2025), the data corruption process is modified to
@@ 4 Experimental Results
Finally, we evaluate both the data-free and adversarial variants on additional flow-matching models, adapting the codebase to their architectural specifics. Notably, only minimal hyperparameter tuning and model-specific customization are required, as summarized in Tables 3 and 4. Implementation details will be available in the forthcoming code release.
@@ 4.1 Understanding the Role of Loss Reweighting
We first examine three extreme forms of loss reweighting, where the generator loss is restricted to one of three disjoint intervals: $t\in(0,\tfrac{1}{3})$, $t\in(\tfrac{1}{3},\tfrac{2}{3})$, or $t\in(\tfrac{2}{3},1)$. Qualitative generations are shown in Figure 3(a–c). We find that restricting to $t\in(\tfrac{2}{3},1)$ is sufficient to produce visually appealing images, though these often lack high-frequency details and diversity. In contrast, $t\in(\tfrac{1}{3},\tfrac{2}{3})$ yields finer detail but with a duller, hazier appearance, while restricting to $t\in(0,\tfrac{1}{3})$ fails to produce reasonable generations. We then consider less extreme reweighting with partially overlapping intervals: $t\in(0,\tfrac{1}{2})$, $t\in(\tfrac{1}{4},\tfrac{3}{4})$, and $t\in(\tfrac{1}{2},1)$. The corresponding qualitative results are shown in Figure 3(d–f). Similar trends are observed, though the effects are less pronounced.
These empirical findings provide intuition for designing $p(t)$ and $w_{t}$. Since the effective timestep distribution $\pi(t)$ depends only on their product, adjusting both is not strictly necessary to preserve the loss structure. In this paper, we fix $p(t)=\mathrm{Logit}\,\mathcal{N}(t;\ln 2,1.6^{2})$ to match the schedule used for finetuning the SANA-Sprint teacher. We set $w_{t}=1-t$. The resulting weight-normalized distribution $\pi(t)$ is shown in the first column, third row of Figure 2. While a systematic study of how varying $p(t)$ and $w_{t}$ affects performance is beyond the scope of this paper, our observation is consistent with Figure 3: stronger emphasis on larger $t$ values (heavier noise) produces visually appealing but less detailed images, and smaller $t$ highlights fine-grained detail at the cost of vividness. Overall, the chosen combination of $p(t)$ and $w_{t}$ yields a $\pi(t)$ with full coverage over $t$, which we find to perform well across all T2I flow-matching models tested in this paper.
@@ 4.2 Distillation of Flow-Matching-Based SANA Models
Model #Steps Params (B) FID CLIP GenEval
@@
SANA Sprint (TrigFlow) 4 0.6 26.97 0.330 0.76
SiD-DiT (SANA, TrigFlow) 4 0.6 25.34 0.331 0.76
@@
SANA Sprint (TrigFlow) 1 1.6 24.60 0.335 0.77
SiD-DiT (SANA, TrigFlow) 4 1.6 23.81 0.336 0.77
@@
Model #Steps Params (B) FID CLIP GenEval
@@
SD3-Medium (base) 28 2.0 24.40 0.336 0.70
Flash SD3 (Chadebec et al., 2025) 4 2.0 25.20 0.328 0.67
SiD-DiT (SD3-Medium) 4 2.0 25.19 0.333 0.70
@@
SD3.5-Medium (base) 40 2.5 21.25 0.340 0.63
SD3.5-Turbo-Medium 8 2.5 21.15 0.337 0.62
SiD-DiT (SD3.5-Medium) 4 2.5 21.07 0.341 0.64
@@
SD3.5-Large (base) 28 8.1 20.81 0.341 0.69
SD3.5-Turbo-Large 4 8.1 26.11 0.340 0.69
SiD-DiT (SD3.5-Large) 4 8.1 20.57 0.341 0.71
FLUX.1-Dev 28 12.0 22.92 0.344 0.65
SiD-DiT (FLUX.1-Dev) 4 12.0 27.86 0.330 0.60
Quantitative results for both rectified-flow- and TrigFlow-based SANA are reported in Table 2. We evaluate performance on the SANA backbone using zero-shot FID, CLIP score (Radford et al., 2021), and GenEval (Ghosh et al., 2023), with FID and CLIP computed on the 10k COCO-2014 validation subset employed by DMD2 (Yin et al., 2024a). While these metrics provide valuable insights for comparing a teacher model with its distilled variants, we caution against using them to directly compare models of different families and sizes. Although SANA-1.6B and its distilled variants may appear to outperform SD3, SD3.5, and FLUX due to higher GenEval scores and comparable FID/CLIP values—as shown in the tables below—visual inspection does not clearly support this conclusion. These metrics remain informative, but their interpretation across families and scales should be made with caution and without overgeneralization.
@@
We evaluate SiD-DiT on SD3-Medium (2B parameters) and SD3.5-Medium (2.5B parameters), both based on the MMDiT architecture (Esser et al., 2024), which improves visual fidelity, typography, complex prompt comprehension, and computational efficiency. Using the same teacher noise schedule as SANA and the $w_{t}=1-t$ reweighting, we observe consistent success across both models, with results summarized in Table 2.
@@
Building on these successes, we extend SiD-DiT to SD3.5-Large (8.1B parameters), the largest open-source MMDiT model currently available in the Stable Diffusion family and more than three times larger than SD3.5-Medium (2.5B). Scaling to this size introduces substantial memory challenges; however, our FSDP+BF16+streaming strategy alleviates these constraints, enabling distillation on a single $\times8$0GB A100/H100 node without CPU offloading. As shown in Table 2, SiD-DiT achieves an FID of 20.57, substantially outperforming SD3.5-Turbo-Large (26.11) and slightly surpassing the teacher baseline (20.81). Its CLIP score (0.341) matches that of the teacher. Additionally, SiD-DiT outperforms the SD3.5-Large teacher model with 0.71 GenEval score. These results demonstrate that SiD-DiT scales effectively to large MMDiTs, providing a practical, out-of-the-box solution for distilling models at this scale.
@@
The SiD-DiT framework delivers competitive generation quality and provides an out-of-the-box DiT distillation method that is robust across diverse architectures. In our implementation, SiD-DiT employs classifier-free guidance (CFG), as formalized in Equation 15, consistent with both the Stable Diffusion and SANA T2I families. In contrast, FLUX.1-DEV adopts a learned guidance embedding by default and does not provide an explicit unconditional branch for CFG. We partially attribute the modest performance gap of SiD-DiT on FLUX.1-DEV to this guidance-mechanism mismatch. Importantly, we did not introduce any Flux-specific modifications beyond the minimal adjustments required to make the model runnable. Even under this direct application, SiD-DiT achieves strong qualitative results (Figure 4) and competitive quantitative metrics (Table 2), while efficiently distilling the 12B-parameter FLUX.1-DEV model at $512\times 512$ resolution on a single node with eight 80GB GPUs, and at $1024\times 1024$ resolution on a single node with eight 192GB GPUs. Further improvements are likely possible by tailoring SiD-DiT more closely to the unique design of FLUX.1-DEV, for example by integrating its learned guidance embeddings into the distillation objective or developing a hybrid approach that blends CFG with model-specific guidance. Such targeted extensions may help close the remaining performance gap and demonstrate the flexibility of SiD-DiT across emerging flow-matching architectures. We leave such extensions to future work, once abundant large-memory computing resources become available.

### v2

@@ Abstract
Diffusion models achieve high-quality image generation but are limited by slow iterative sampling. Distillation methods alleviate this by enabling one- or few-step generation. Flow matching, originally introduced as a distinct framework, has since been shown to be theoretically equivalent to diffusion under Gaussian assumptions, raising the question of whether distillation techniques such as score distillation transfer directly. We provide a simple derivation—based on Bayes’ rule and conditional expectations—that unifies Gaussian diffusion and flow matching without relying on ODE/SDE formulations. Building on this view, we extend Score identity Distillation (SiD) to pretrained text-to-image flow-matching models, including SANA, SD3-Medium, SD3.5-Medium/Large, and FLUX.1-dev, all with DiT backbones. Experiments show that, with only modest flow-matching- and DiT-specific adjustments, SiD works out of the box across these models, in both data-free and data-guided settings, without requiring teacher finetuning or architectural changes. This provides the first systematic evidence that score distillation applies broadly to text-to-image flow matching models, resolving prior concerns about stability and soundness and unifying acceleration techniques across diffusion- and flow-based generators. A project page is available at https://yigu1008.github.io/SiD-DiT.
Figure 1: Qualitative results produced by the four-step SiD-DiT generator distilled from SD3.5-Large.
@@ 1 Introduction
This theoretical equivalence raises a natural question: can diffusion distillation techniques—broadly divided into trajectory and score distillation (Fan et al., 2025), and proven effective for compressing pretrained diffusion models into one- or few-step generators—be directly applied to flow-matching models? Prior work has begun to explore this. The continuous-time consistency model (Lu & Song, 2024) introduced TrigFlow and demonstrated trajectory distillation for pretrained TrigFlow models. Extending this to text-to-image (T2I) generation, Chen et al. (2025) developed SANA-Sprint by reformulating SANA (Xie et al., 2024) from rectified flow into TrigFlow and applying consistency distillation. While effective, this approach requires nontrivial finetuning of rectified-flow checkpoints into TrigFlow counterparts, making it inapplicable to pretrained rectified-flow models without additional adaptation.
Score distillation relaxes the constraint of strictly following the teacher’s sampling trajectory and has shown consistent gains over trajectory-based consistency distillation on diffusion benchmarks such as CIFAR-10 and ImageNet (Zhou et al., 2025c). Yet its applicability to flow-matching T2I models remains unclear. If effective, a further question is whether additional adaptation steps—such as finetuning, as in SANA-Sprint—are necessary. This uncertainty is compounded by a sensitive design space, including noise schedules, loss weighting, network preconditioning (Karras et al., 2022), and architecture. Small changes in these factors can significantly affect performance, as evidenced by methods like sCM, which require careful adaptation during pretraining (Lu & Song, 2024) or finetuning (Chen et al., 2025). Concerns about stability further complicate matters: consistency distillation was favored in SANA-Sprint partly due to instability observed in Distribution Matching Distillation (DMD) (Yin et al., 2024c; a). However, it remains unclear whether this instability is unique to DMD’s KL-based formulation or reflects broader issues in score distillation, which can also be defined with divergences such as Fisher divergence (Zhou et al., 2024) or $f$-divergences (Xu et al., 2025). Huang et al. (2024) argue flow matching does not explicitly model probability density, raising doubts about the soundness of applying distribution-divergence-based objectives directly.
@@
To validate this view, we adopt the few-step Score identity Distillation (SiD) framework (Zhou et al., 2025a), previously shown effective for diffusion models such as SD1.5 and SDXL with U-Net backbones. Here, we extend SiD to pretrained flow-matching models with Diffusion Transformer (DiT) (Peebles & Xie, 2023) backbones, including SANA (Xie et al., 2024; Chen et al., 2025), SD3-Medium, SD3.5-Medium, SD3.5-Large, and FLUX.1-dev (Labs, 2024), spanning 0.6B–12B parameters (2.4–48 GB in fp32). We show that SiD works out of the box across these models in both data-free and data-guided settings: the former requires no additional images beyond the teacher, while the latter incorporates adversarial learning by pooling discriminator features along the spatial dimension from a suitable DiT layer without introducing new parameters.
We provide a review of related work in Appendix B. Code and additional results are available at our project page: https://yigu1008.github.io/SiD-DiT. Importantly, a single codebase and hyperparameter configuration suffice across all T2I flow-matching models, underscoring the robustness and applicability of the SiD-DiT framework.
@@ 2.2 Equivalence of Diffusion and Flow-Matching Objectives and Variants
TrigFlow. In TrigFlow (Lu & Song, 2024), the data corruption process is modified to
@@ 4 Experimental Results
Finally, we evaluate both the data-free and adversarial variants on additional flow-matching models, adapting the codebase to their architectural specifics. Notably, only minimal hyperparameter tuning and model-specific customization are required, as summarized in Tables 4 and 5. As shown in Figure 4, SiD-DiT achieves rapid improvements in both FID and CLIP scores during distillation across all nine DiT models. Full implementation details are provided in the supplementary code release.
@@ 4.1 Understanding the Role of Loss Reweighting
We first examine three extreme forms of loss reweighting, where the generator loss is restricted to one of three disjoint intervals: $t\in(0,\tfrac{1}{3})$, $t\in(\tfrac{1}{3},\tfrac{2}{3})$, or $t\in(\tfrac{2}{3},1)$. Qualitative generations are shown in Figure 3(a–c).
Figure 4: This plot shows the evolution of FID (solid lines, left y-axis) and CLIP score (matching line styles with reduced opacity, right y-axis) as a function of the number of iterated images (in thousands) for SiD-DiT. Because the x-axis is log-scaled, the near-linear trends in many panels reflect a rapid initial decline in FID accompanied by a corresponding rise in CLIP score, followed by progressively smaller gains as training continues. This consistent behavior across architectures and model sizes shows that SiD-DiT quickly improves both image fidelity and semantic alignment during the early stages of distillation.
We find that restricting to $t\in(\tfrac{2}{3},1)$ is sufficient to produce visually appealing images, though these often lack high-frequency details and diversity. In contrast, $t\in(\tfrac{1}{3},\tfrac{2}{3})$ yields finer detail but with a duller, hazier appearance, while restricting to $t\in(0,\tfrac{1}{3})$ fails to produce reasonable generations. We then consider less extreme reweighting with partially overlapping intervals: $t\in(0,\tfrac{1}{2})$, $t\in(\tfrac{1}{4},\tfrac{3}{4})$, and $t\in(\tfrac{1}{2},1)$. The corresponding qualitative results are shown in Figure 3(d–f). Similar trends are observed, though the effects are less pronounced.
These empirical findings provide intuition for designing $p(t)$ and $w_{t}$. Since the effective timestep distribution $\pi(t)$ depends only on their product, adjusting both is not strictly necessary to preserve the loss structure. In this paper, we fix $p(t)=\mathrm{Logit}\,\mathcal{N}(t;\ln 2,1.6^{2})$ to match the schedule used for finetuning the SANA-Sprint teacher. We set $w_{t}=1-t$. The resulting weight-normalized distribution $\pi(t)$ is shown in the first column, third row of Figure 2. While a systematic study of how varying $p(t)$ and $w(t)$ affects performance is beyond the scope of this paper, our observation is consistent with Figure 2: stronger emphasis on larger $t$ values (heavier noise) produces visually appealing but less detailed images, and smaller $t$ highlights fine-grained detail at the cost of vividness. Overall, the chosen combination of $p(t)$ and $w_{t}$ yields a $\pi(t)$ with full coverage over $t$, which we find to perform well across all T2I flow-matching models tested in this paper.
@@ 4.2 Distillation of Flow-Matching-Based SANA Models
Model #Steps Params (B) FID $\downarrow$ CLIP $\uparrow$ GenEval $\uparrow$ Aesth. $\uparrow$ HPSv2 $\uparrow$ ImgRwd $\uparrow$ PickScore $\uparrow$
@@
SANA Sprint Teacher 20 0.6 25.64 0.335 0.780 6.222 0.299 1.115 21.78
SANA Sprint (Chen et al., 2025) (TrigFlow, 1 step) 1 0.6 24.60 0.336 0.770 6.361 0.286 1.006 21.805
SANA Sprint (Chen et al., 2025) (TrigFlow, 4 steps) 4 0.6 26.32 0.335 0.766 6.325 0.301 1.111 22.125
SiD-DiT (SANA, TrigFlow) 4 0.6 25.81 0.340 0.763 6.243 0.308 1.049 21.561
@@
SANA Sprint Teacher 20 1.6 25.64 0.335 0.776 6.209 0.304 1.163 21.93
SANA Sprint (Chen et al., 2025) (TrigFlow, 1 step) 1 1.6 24.60 0.335 0.768 6.362 0.293 1.030 22.006
SANA Sprint (Chen et al., 2025) (TrigFlow, 4 steps) 4 1.6 24.79 0.335 0.768 6.338 0.300 1.081 22.123
SiD-DiT (SANA, TrigFlow) 4 1.6 23.81 0.340 0.774 6.305 0.307 1.102 21.897
@@
Model #Steps Params (B) FID $\downarrow$ CLIP $\uparrow$ Aesth. $\uparrow$ HPSv2 $\uparrow$ ImgRwd $\uparrow$ PickScore $\uparrow$
@@
SD3-Medium (base) 28 2.0 24.40 0.336 5.870 0.297 1.051 21.574
Flash SD3 (Chadebec et al., 2025) 4 2.0 22.70 0.338 5.820 0.289 0.997 21.326
SiD-DiT (SD3-Medium) 4 2.0 22.05 0.341 6.054 0.301 1.017 21.686
@@
SD3.5-Medium (base) 40 2.5 22.51 0.342 5.973 0.300 1.089 21.974
SD3.5-Medium-Turbo 8 2.5 21.15 0.337 5.971 0.263 0.633 21.478
SiD-DiT (SD3.5-Medium) 4 2.5 21.07 0.340 6.187 0.308 1.097 22.037
@@
SD3.5-Large (base) 28 8.1 20.81 0.341 6.097 0.305 1.127 22.245
SD3.5-Turbo-Large 4 8.1 26.11 0.340 6.198 0.302 1.086 22.171
SiD-DiT (SD3.5-Large) 4 8.1 21.10 0.341 6.132 0.309 1.214 22.097
[…]
FLUX-1 Family
FLUX-1-Dev (base) 28 12.0 22.89 0.344 6.184 0.297 0.897 21.862
Flux-Schnell 4 12.0 23.42 0.345 6.173 0.302 1.109 21.796
FLUX-1-Turbo 4 12.0 24.92 0.332 6.192 0.302 1.012 21.977
Hyper-FLUX 4 12.0 25.44 0.332 6.257 0.310 1.028 22.090
SiD-DiT (FLUX-1-Dev) 4 12.0 27.86 0.330 5.964 0.305 1.203 21.583
Quantitative results for both rectified-flow- and TrigFlow-based SANA are reported in Table LABEL:tab:sana-quantitative. We evaluate performance on the SANA backbone using zero-shot FID, CLIP score (Radford et al., 2021), and GenEval (Ghosh et al., 2023), with FID and CLIP computed on the 10k COCO-2014 validation subset employed by DMD2 (Yin et al., 2024a). We also evaluate human preference using LAION Aesthetics (Schuhmann et al., 2021), HPSv2 (Wu et al., 2023), ImageReward (Xu et al., 2023), and PickScore  (Kirstain et al., 2023) on 2048 Pick-a-Pic  (Kirstain et al., 2023) validation prompts.
@@
We evaluate SiD-DiT on SD3-Medium (2B parameters) and SD3.5-Medium (2.5B parameters), both based on the MMDiT architecture (Esser et al., 2024), which improves visual fidelity, typography, complex prompt comprehension, and computational efficiency. Using the same teacher noise schedule as SANA and the $w_{t}=1-t$ reweighting, we observe consistent success across both models, with results summarized in Table LABEL:tab:sd3-quantitative.
@@
[…]
@@
The SiD-DiT framework delivers competitive generation quality and serves as an out-of-the-box DiT distillation method that is robust across diverse architectures. In our implementation, SiD-DiT employs CFG as formalized in Equation 15, consistent with the Stable Diffusion T2I family. In contrast, FLUX.1-DEV adopts a learned guidance embedding by default and does not provide an explicit unconditional branch for CFG. We partially attribute the modest performance gap of SiD-DiT on FLUX.1-DEV to this guidance-mechanism mismatch. Importantly, we did not introduce any Flux-specific modifications beyond the minimal adjustments required to make the model runnable. Even under this direct application, SiD-DiT achieves strong qualitative results (Figure 5) and competitive quantitative metrics (Table LABEL:tab:sd3-quantitative), while efficiently distilling the 12B-parameter FLUX.1-DEV model at $512\times 512$ resolution on a single node with eight 80GB GPUs, and at $1024\times 1024$ resolution on a single node with eight 192GB GPUs. Further improvements are likely possible by tailoring SiD-DiT more closely to the unique design of FLUX.1-DEV, for example by integrating its learned guidance embeddings into the distillation objective or developing a hybrid approach that blends CFG with model-specific guidance. Such targeted extensions may help close the remaining performance gap and demonstrate the flexibility of SiD-DiT across emerging flow-matching architectures.